# SIR: Multi-view Inverse Rendering with Decomposable Shadow for Indoor Scenes

### 25 Feb 2024 v2

@@ 3.4 BRDF
where $\Phi(\cdot)$ is an operator to obtain the V value in HSV color domain by converting RGB-space to HSV-space, $\Phi_{inv}(\cdot)$ is an inverse operator to convert HSV-space to RGB-space [20], $M_{i}$ is the prior instance segmentation mask, $K$ denotes the instance classes of the indoor scene, $p$ denotes the minibatch of 3D points, and $\odot$ denotes an element-wise product. Meanwhile, we introduce a roughness smooth regularization $\mathcal{L}_{rough}$, which constrains the roughness value on the same instance level to be similar.
@@
3.5 Training
where $\mathcal{X}$ is the minibatch of 3D points uniformly sampled in 3D space and nearby surface. $\mathcal{L}_{normal}$ is a normal smooth loss [45], which constrains the predicted normals $\bm{n}_{i}$ connected to the sampled density gradient normals $\bar{\bm{n}}_{i}$ along the ray. Formally:
Phase 2: Training of hard shadow and irradiance field. In Sec. 3.2 and Sec. 3.3, we can obtain the incoming diffuse lighting $L_{i,d}$ and hard shadow $\xi$, which serve as the ground truth for supervising the predicted irradiance $I_{ir}$ and hard shadow $S_{hard}$ via $\ell_{1}$ loss.
$\mathcal{L}_{2}=||L_{i,d}(\hat{\textbf{x}})-{I_{ir}}(\hat{\textbf{x}})||_{1}+||\xi(\hat{\textbf{x}})-S_{hard}(\hat{\textbf{x}})||_{1}$ (19)
Phase 3: Training of BRDF and soft shadow field. During the optimization of intrinsic decomposition, we adopt a multi-stage training scheme to train BRDF and soft shadow field to avoid the ambiguity between albedo and shadow. We first train the initial albedo under shadow and roughness layer parameters fixed, then we train the albedo and shadow with $\mathcal{L}_{albedo}$ under roughness layer parameters fixed, and we finetune the roughness with $\mathcal{L}_{rough}$ after the albedo and shadow MLP have been trained.
where $\mathcal{L}_{render}$ is = $||\hat{L_{o}}(\hat{\textbf{x}})-L_{o}(\hat{\textbf{x}})||_{1}$, minimizing the $\ell_{1}$ error between the re-rendered (see Eq. 9) and GT images. More details of the rendering equation can refer to Sec. A and  B.1 in the Appendix.
4 Experiments
Table 2: Quantitative results for all baselines and ours. All scores are calculated as an average across 2 scenes from the real-world dataset.
View Synthesis
@@ 4 Experiments
4.1 Indoor Datasets
Owing to the absence of shadow effects within the existing HDR multi-view datasets for indoor scenes, we recognize the need for a more robust evaluation of our methodology. Consequently, we have developed two distinct inverse rendering datasets tailored specifically for indoor scene evaluation: 1) The synthetic dataset is an extension of the DM-NeRF dataset [46], featuring six indoor scenes. It includes 300 posed images for training and 200 for testing. Each image has $400\times 400$ pixels. We alter the locations and colors of objects within these scenes and enhance various object properties (e.g., albedo, roughness, shadow, instance mask, etc.). 2) Our real-world dataset includes two indoor scenes with elaborate materials and sophisticated lighting conditions. Utilizing a professional camera, we capture 120 HDR images for each scene, merging each HDR image from three distinct exposures (from $\frac{1}{15000}$ s to $\frac{1}{8}$ s). The captured images are processed to a resolution of $512\times 512$ pixels, 85% for training and 15% for testing. These two datasets provide challenging indoor scenes for inverse rendering analysis.
4.2 Implementation
All neural fields in our network are implemented by multi-layer perceptrons (MLPs) with ReLU activations. For the geometry network $f_{d}$ and HDR-radiance network $f_{c}$, we follow the default setting from VolSDF  [16], where $f_{d}$ is an 8-layer MLP with hidden dimension 256 and $f_{c}$ is a 4-layer MLP with hidden dimension 256. The architectures of irradiance MLP, shadow MLPs, and BRDF MLP all contain 4 layers with 256 hidden units. Positional encoding [5] is applied to the input 3D locations and directions with 10 and 4 frequency components, respectively. For the SDF-based neural radiance field, we implement the two networks in PyTorch and optimize using Adam [47] with a learning rate of $5\times 10^{-4}$ for 250K iterations. In irradiance and hard shadow estimation, we use the Adam optimizer with a learning rate of $5\times 10^{-4}$ for 10K iterations, and we use 512 samples to compute the irradiance (Sec. 3.2) and shadow (Sec. 3.3). In all three stages of material estimation, we use the Adam optimizer with a learning rate of $10^{-3}$ for 25K iterations, and we use 128 samples to compute the BRDF term (Sec. 3.4). All experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU.
@@ 4.2 Implementation
4.3 Inverse Rendering
We select four notable and widely recognized inverse rendering methods as our baselines: 1) NVDIFFREC [12], 2) IBL-NeRF [36], 3) PhySG [32], and 4) InvRender [33]. Specifically, NVDIFFREC employs a DMTet-based approach, utilizing a differentiable marching tetrahedral layer to learn object-level meshes. IBL-NeRF, on the other hand, is a NeRF-based method for inverse rendering. PhySG and InvRender use an SDF-based neural radiance field for inverse rendering, which are both evaluated on object-level datasets. For fair comparisons on indoor scenes, we have minor modifications to adapt these baselines. Specifically, we align the geometric backbone of both InvRender [33] and PhySG [32] with our geometry representation (i.e., VolSDF), considering that both use SDF-based neural radiance fields for geometry. Further details are provided in Sec. B.2 in Appendix.
We report image quality metrics: PSNR, SSIM [48], and LPIPS [49] on test viewpoints for view synthesis results and albedo, as well as the mean squared error (MSE) for roughness and shadow predictions. The evaluations of our method, as evidenced by the quantitative results in Table 1 and qualitative results in Fig. 3, demonstrate that our SIR outperforms existing methods in decomposing material attributes like albedo and roughness, and more notably, in shadow extraction. This leads to superior view synthesis results compared to other methods on indoor scenes. These results convincingly demonstrate the efficiency of the proposed shadow term in enhancing the overall view synthesis. Furthermore, as shown in Table 2 and Fig. 3, our method exhibits a remarkable capability to handle complex lighting conditions in real-world indoor scenes. This ability is crucial for achieving precise and reliable material estimations for inverse rendering.
4.4 Ablation Study
The key innovation of SIR is the capacity to effectively isolate shadows while ensuring that these shadows are not incorporated into the albedo under an indoor scene. Therefore, this section is dedicated to evaluating the impact of shadow terms, differentiable soft shadow, and albedo regularizer:
(1) Removing shadow terms: Without the shadow terms means removing both $S_{hard}$ in Eq. 12 and $S_{soft}$ in Eq. 13, as well as remove $S_{hard}$ in the training of phase 2 (see Eq. 19). This is to demonstrate the necessity of shadow terms in the rendering equation for accurate inverse rendering.
(2) Removing soft shadow: Without using the differentiable soft shadow $S_{soft}$ in phase 3, the pipeline re-renders images from hard shadows $S_{hard}$ learned in phase 2.
(3) Removing albedo regularizer: This ablation is to evaluate the enhancement of introducing instance-level albedo consistency as a regularizer.
Table 3 (1) clearly shows that removing the shadow term in the rendering equation, leads to a remarkable decline in the performance of novel view synthesis, adversely affecting the albedo estimation as well. The qualitative results in Fig. 4 further indicate that the absence of shadow terms results in incorrect albedo outputs with baked shadows, which then negatively impact the synthesis results. However, intuitively introducing a binary (hard) shadow still leads to suboptimal synthesis as shown in Table 3 (2), particularly noticeable in the inaccuracies around shadow edges (see Fig. 4). Therefore, learning a differentiable soft shadow after an initial hard shadow is necessary to preserve the shadow details in the synthesis images. Moreover, as demonstrated in Table 3 (3), the integration of an instance-level albedo regularizer proves to be effective in maintaining consistency in albedo for each instance, thereby contributing to more accurate synthesis results.
@@ 4.5 Scene Editing
In Fig. 5, we present the results of editing six varied indoor scenes, encompassing both synthetic and real-world scenarios. The initial row displays the synthesized view images from our test set, serving as the source images for subsequent edits. The second row illustrates the outputs of free-viewpoint relighting experiments, featuring varied light positions and colors. Thanks to the cleanly decomposed albedo, free from any pre-existing shadows, our relighting experiments exhibit new realistic distributions of shadows. The third row is dedicated to object insertion, where we have incorporated six models from the Stanford 3D models [50] into these scenes in diverse positions with different sizes and colors to compare the effect on re-rendering and shadow calculation. Our method achieves realistic shadowing and shading effects for these insertions. Furthermore, the final row shows material replacement: we randomly altered materials within the scenes, experimenting with different textures and levels of roughness (e.g., the wall in the Bathroom scene). These convincingly realistic editing results demonstrate the effectiveness and robustness of our SIR in decomposing the intrinsic properties of indoor scenes, which is fundamental for accurate inverse rendering applications.
@@ 5 Conclusion
We propose SIR, a multi-view inverse rendering method on indoor scenes, effectively addressing the challenges of material and lighting decomposition by explicitly isolating shadows. Leveraging posed HDR images and an SDF-based neural radiance field, SIR significantly enhances realism in material estimation and scene editing, surpassing previous methods with accurate shadow estimation. The innovative incorporation of a hard shadow term within a three-stage material estimation process substantially improves the quality of SVBRDFs. Extensive evaluations on diverse indoor datasets demonstrate the superiority of SIR in both quantitative and qualitative aspects over existing methods. Note that, since our research primarily addresses the challenge of decomposing shadows and albedo, we directly utilize existing methods for geometric representation. The effectiveness of our approach depends partly on the accuracy of geometric inputs. Therefore, incorporating more precise geometric representations could further improve inverse rendering quality.
Acknowledgement
This research is funded by the Laboratory for Artificial Intelligence in Design (Project Code: RP1-3), Innovation and Technology Fund, Hong Kong Special Administrative Region, and PolyU project P0040352.
Appendix A BRDF Model
In Sec 3.4, we adapt the Microfacet BRDF model [51, 43] to approximate the surface reflectance property with a set of decomposed intrinsic terms, which can encapsulate the interaction of light with a surface, and the microfacet model provides a robust framework for simulating the reflection from rough surfaces. The BRDF $f_{r}(\hat{\textbf{x}},\bm{\omega}_{i},\bm{\omega}_{o})$ is the ratio between the incoming and outgoing radiance and it is a function of the surface location $\hat{\textbf{x}}$, incoming direction $\bm{\omega}_{i}$, and the outgoing direction $\bm{\omega}_{o}$.
@@ Appendix A BRDF Model
where $\hat{\bm{n}}$ is the surface normal at $\hat{\textbf{x}}$. The BRDF is composed of two terms diffuse reflection $f_{d}$ and specular reflection $f_{s}$. In the $1$th stage of BRDF estimation, we use the diffuse reflection term $f_{d}(\hat{\textbf{x}})=\frac{\hat{A}_{x}}{\pi}$ based on a simple Lambertian model. Here, we introduce the specular BRDF $f_{s}$ which follows the Cook-Torrance model [52]:
@@
where $\hat{R}$ is roughness term, $\bm{h}$ is a half vector, $D$ denotes Normal Distribution Function (NDF), $F$ denotes Fresnel function and $G$ is the Geometry Factor.
@@
$D(h)=\frac{\alpha^{2}}{\pi((\hat{\bm{n}}\cdot\bm{h})^{2}(\alpha^{2}-1)+1)^{2}}$ (23)
where the half-vector $\bm{h}$ is the normalized vector halfway between the incident light direction and the reflection direction, $\hat{\bm{n}}$ is the macroscopic surface normal, and $\alpha$ is the surface roughness parameter, ranging from [0,1]. As $\alpha$ approaches 0, the surface tends towards a mirror-like reflection, while $\alpha$ approaching 1 indicates an extremely rough surface.
@@
$F(\bm{\omega}_{o},\bm{h})=F_{0}+(1-F_{0})\cdot(1-(\hat{\bm{n}}\cdot\bm{h}))^{5}$ (24)
where $F_{0}$ is the reflectance at normal incidence, also known as the base reflectance and $\hat{\bm{n}}\cdot\bm{h}$ is the dot product between the surface normal and the half-vector, dictating the cosine of the angle of incidence. In our method, we assume dielectric materials with $F_{0}=0.04$ in the Fresnel term.
@@
Appendix B Additional Implementation Details
B.1 Implemetaions of SIR
In Sec. 3.5, we use a three-phase training strategy to implement our overall framework, which includes the geometry and HDR radiance field, irradiance and hard shadow field, and material and soft shadow estimation. The details are as follows:

### 09 Apr 2024 v3

@@ 3.5 BRDF
3.6 Training
where $\mathcal{X}$ is the minibatch of 3D points uniformly sampled in 3D space and nearby surface. $\mathcal{L}_{normal}$ is a normal smooth loss [44], which constrains the predicted normals $\bm{n}_{i}$ connected to the sampled density gradient normals $\bar{\bm{n}}_{i}$ along the ray:
Phase 2: Training of hard shadow and irradiance field. In Sec. 3.3 and Sec. 3.4, we can obtain the incoming diffuse lighting $L_{i,d}$ and hard shadow $\xi$, which serve as the ground truth for supervising the predicted irradiance $I$ and hard shadow $S_{hard}$ via $\ell_{1}$ loss.
$\mathcal{L}_{2}=||L_{i,d}-{I}||_{1}+||\xi-S_{hard}||_{1}$ (15)
Phase 3: Training of BRDF and soft shadow field. During the optimization of intrinsic decomposition, we adopt a three-stage training scheme in Sec. 3.5 to train BRDF and soft shadow field to avoid ambiguities between albedo and shadow. In stage 1, we optimize $\mathcal{L}_{render}$ = $||L_{o}-\hat{L}_{o}||_{1}$ to update the coarse albedo $\hat{A}$ with fixed hard shadow $S_{hard}$ and irradiance $I$. Then, in stage 2, we learn the differentiable soft shadow $S_{soft}$ and refine albedo by Eq. 16 with $\lambda_{albedo}=10^{-4}$ and $\lambda_{rough}=0$. Finally, we fix the above parameters in diffuse term and refine roughness values by Eq. 16 with $\lambda_{albedo}=0$ and $\lambda_{rough}=10^{-4}$.
More details of the rendering equation and implementation can refer to Sec. 0.A and  0.B.1 in the Appendix.
4 Experiments
4.1 Indoor Datasets
Owing to the absence of shadow effects within the existing HDR multi-view datasets for indoor scenes, we recognize the need for a more robust evaluation of our methodology. Consequently, we have developed two distinct inverse rendering datasets tailored specifically for indoor scene evaluation: 1) The synthetic dataset is an extension of the DM-NeRF dataset [46], featuring six indoor scenes generated using Blender Cycles Render. It includes 300 posed images for training and 200 for testing. Each image has $400\times 400$ pixels. We alter the locations and colors of objects within these scenes and enhance various object properties (e.g., albedo, roughness, shadow, instance mask, etc.). 2) Our real-world dataset includes two indoor scenes with elaborate materials and sophisticated lighting conditions. Utilizing a professional camera, we capture 120 HDR images for each scene, merging each HDR image from three distinct exposures (from $\frac{1}{15000}$ s to $\frac{1}{8}$ s). The captured images are processed to a resolution of $512\times 512$ pixels, 85% for training and 15% for testing. These two datasets provide challenging indoor scenes for inverse rendering analysis.
4.2 Implementation
All neural fields in our network are implemented by multi-layer perceptrons (MLPs) with ReLU activations. For the geometry network $f_{d}$ and HDR-radiance network $f_{c}$, we follow the default setting from VolSDF [49], where $f_{d}$ is an 8-layer MLP with hidden dimension 256 and $f_{c}$ is a 4-layer MLP with hidden dimension 256. The architectures of irradiance MLP, shadow MLPs, and BRDF MLP all contain 4 layers with 256 hidden units. Positional encoding [34] is applied to the input 3D locations and directions with 10 and 4 frequency components, respectively. For the SDF-based neural radiance field, we implement the two networks in PyTorch and optimize using Adam [21] with a learning rate of $5\times 10^{-4}$ for 250K iterations. For irradiance and hard shadow estimation, we use the Adam optimizer with a learning rate of $5\times 10^{-4}$ for 10K iterations, and we use 512 sample rays at any surface point to compute the irradiance (Sec. 3.3) and shadow (Sec. 3.4). In all three stages of material estimation, we use the Adam optimizer with a learning rate of $10^{-3}$ for 25K iterations, and we use 128 samples to compute the BRDF term (Sec. 3.5). All experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU. The entire training process takes approximately 40 hours and is divided into three phases: 16 hours for the first phase, 8 hours for the second phase, and 16 hours for the third phase.
4.3 Inverse Rendering
We select four notable and widely recognized inverse rendering methods as our baselines: 1) NVDIFFREC [36], 2) IBL-NeRF [4], 3) PhySG [53], and 4) InvRender [57]. Specifically, NVDIFFREC employs a DMTet-based approach, utilizing a differentiable marching tetrahedral layer to learn object-level meshes. IBL-NeRF, on the other hand, is a NeRF-based method for inverse rendering. PhySG and InvRender use an SDF-based neural radiance field for inverse rendering, which are both evaluated on object-level datasets. For fair comparisons on indoor scenes, we have minor modifications to adapt these baselines. Specifically, we align the geometric backbone of both InvRender [57] and PhySG [53] with our geometry representation (i.e., VolSDF [49]), considering that both use SDF-based neural radiance fields for geometry. Further details are provided in Sec. 0.B.2 in Appendix.
For the synthetic dataset, we report image quality metrics: PSNR, SSIM [18], and LPIPS [55] on test viewpoints for view synthesis results and albedo, as well as the mean squared error (MSE) for roughness and shadow predictions. In particular, since our SIR is the unique method that individually recovers shadows, we can only report the MSE on shadow predictions of our method. Besides, the ground truth shadows generated by the Blender renderer are binary, so we convert our soft shadows to binary for metric computation. The evaluations of our method, as evidenced by the quantitative results in Table 1 and qualitative results in Fig. 3, demonstrate that our SIR outperforms existing methods in decomposing material attributes like albedo and roughness, and more notably, in shadow extraction. This leads to superior view synthesis results compared to other methods on indoor scenes. These results convincingly demonstrate the efficiency of the proposed shadow term in enhancing the overall view synthesis.
Furthermore, we report view synthesis results of the real-world dataset. Our method exhibits a remarkable capability to handle complex lighting conditions in real-world indoor scenes in Table 3 and Fig. 3, which is crucial for achieving reliable material estimations for inverse rendering. The PSNR of our view synthesis results are higher than most of the baselines and slightly lower than IBL-NeRF. Meanwhile, the values of SSIM and LPIPS indicate that our results have less distortion and greater perceptual similarity to the targets.
@@ 4.3 Inverse Rendering
Table 2: Quantitative results on relighting for all baselines and ours. All scores are calculated as an average across 6 scenes from the synthetic dataset.
Table 3: Quantitative results on view synthesis for all baselines and ours. All scores are calculated as an average across 2 scenes from the real-world dataset.
[…]
NVDIFFREC [36] 14.1443 0.7175 0.5041
PhySG [53] 13.5117 0.7716 0.2385
InvRender [57] 14.3757 0.7240 0.3746
SIR (Ours) 21.4497 0.8215 0.2941
@@
Table 3: Quantitative results on view synthesis for all baselines and ours. All scores are calculated as an average across 2 scenes from the real-world dataset.
Figure 3: Qualitative results of all methods on two datasets (Syn: Restroom, Real: Office).
4.4 Ablation Study
The key innovation of our method is the capacity to effectively isolate shadows while ensuring that these shadows are not incorporated into the albedo under an indoor scene. Therefore, this section is dedicated to evaluating the impact of shadow terms, differentiable soft shadow, and albedo regularizer:
(1) Removing shadow terms: We remove both $S_{hard}$ and $S_{soft}$ in phases 2, 3 to demonstrate the necessity of introducing shadow terms in inverse rendering.
(2) Removing soft shadow: Without the differentiable shadow $S_{soft}$ in phase 3, the pipeline re-renders images using hard shadows $S_{hard}$.
(3) Removing albedo regularizer: This ablation is to evaluate the enhancement of introducing instance-level albedo consistency.
Table 4 (1) clearly shows that removing the shadow term in the rendering equation, leads to a remarkable decline in the performance of novel view synthesis, adversely affecting the albedo estimation as well. The qualitative results in Fig. 4 further indicate that the absence of shadow terms results in incorrect albedo outputs with baked shadows, which then negatively impact the synthesis results. However, intuitively introducing a binary (hard) shadow still leads to suboptimal synthesis as shown in Table 4 (2), particularly in the inaccuracies around shadow edges (see Fig. 4). Therefore, learning a differentiable soft shadow after an initial hard shadow is necessary to preserve the shadow details in the synthesis images. Moreover, as demonstrated in Table 4 (3), the integration of an instance-level albedo regularizer proves to be effective in maintaining consistency in albedo for each instance, thereby contributing to more accurate synthesis results. Furthermore, the MSE values on roughness in Table 4 clearly show that correctly decomposing shadows and albedo can prevent the tendency towards non-convergence in roughness, so as to recover accurate roughness values.
@@ 4.4 Ablation Study
Figure 4: Qualitative results of ablation study (Kitchen in synthetic dataset).
Figure 5: Qualitative results of scene editing using our SIR on synthetic dataset.
@@ 4.5 Scene Editing
For quantitative evaluation on relighting of the synthetic dataset, we report the image quality metrics on test viewpoints as shown in Table 3. Due to the limited capability of the irradiance module of IBL-NeRF in relighting, we can only compare the relighting performances with the other three baselines. Although other baselines have attempted to regularize albedo, they fail to account for the shadow term, which means albedo remains entangled with most of the shadows, significantly degrading the relighting quality. InvRender [57], which accounts for visibility due to self-occlusion, only captures a minor portion of shadows generated by light source occlusion, and thus fails to significantly improve results.
In Fig. 5, we present the results of editing varied indoor scenes, encompassing both synthetic and real-world scenarios. The initial row at left shows the synthesized view images from our test set, serving as the source images for subsequent edits. The second row at left illustrates the outputs of free-viewpoint relighting experiments, featuring varied light positions and colors. Thanks to the cleanly decomposed albedo, free from any pre-existing shadows, our relighting experiments exhibit new realistic distributions of shadows. The first row at right is dedicated to object insertion, where we have incorporated four models from the Stanford 3D models [6] into these scenes in diverse positions with different sizes and colors to compare the effect on re-rendering and shadow calculation. Our method achieves realistic shadowing and shading effects for these insertions. Furthermore, the second row at right shows material replacement: we randomly altered materials within the scenes, experimenting with different textures and levels of roughness (e.g., the wall in the Bathroom scene). These convincingly realistic editing results demonstrate the effectiveness and robustness of our SIR in decomposing the intrinsic properties of indoor scenes, which is fundamental for accurate inverse rendering applications.
@@ 5 Conclusion
We propose SIR, a multi-view inverse rendering method on indoor scenes, effectively addressing the challenges of material and lighting decomposition by explicitly isolating shadows. Leveraging posed HDR images and an SDF-based radiance field, SIR significantly enhances realism in material estimation and scene editing, surpassing previous methods with accurate shadow estimation. The innovative incorporation of shadows with a three-stage material estimation process substantially improves the quality of SVBRDFs. Extensive evaluations on diverse indoor datasets demonstrate the superiority of SIR in both quantitative and qualitative aspects over existing methods.
Limitation. First, our shadow and material estimation rely on fine geometry as an input. Second, for simplification, we neglect to refine the position of light sources for material estimation. In the future, we are interested in expanding our method to address the above limitations.
Appendix 0.A BRDF Model
In Sec 3.5, we adapt the microfacet BRDF model [45, 17] to approximate the surface reflectance property with a set of decomposed intrinsic terms, which can encapsulate the interaction of light with a surface. The microfacet model provides a robust framework for simulating the reflection from rough surfaces. The BRDF $f_{r}(\hat{\textbf{x}},\bm{\omega}_{i},\bm{\omega}_{o})$ is the ratio between the incoming and outgoing radiance. It is a function of the surface location $\hat{\textbf{x}}$, incoming direction $\bm{\omega}_{i}$, and the outgoing direction $\bm{\omega}_{o}$:
@@ Appendix 0.A BRDF Model
where the BRDF comprises diffuse reflection $f_{d}$ and specular reflection $f_{s}$. In the first stage of BRDF estimation, we use the diffuse reflection term $f_{d}(\hat{\textbf{x}})=\frac{\hat{A}_{x}}{\pi}$ based on a simple Lambertian model. Here, we introduce the specular BRDF $f_{s}$ which follows the Cook-Torrance model [5]:
@@
where $\hat{\bm{n}}$ is the surface normal at $\hat{\textbf{x}}$, $\hat{R}$ is roughness term, $\bm{h}$ is a half vector, $D$ denotes Normal Distribution Function (NDF), $F$ denotes Fresnel function and $G$ is the Geometry Factor.
@@
$D(\bm{h},\hat{\bm{n}},\hat{R})=\frac{\alpha^{2}}{\pi((\hat{\bm{n}}\cdot\bm{h})^{2}(\alpha^{2}-1)+1)^{2}}$ (19)
where the half-vector $\bm{h}$ is the normalized vector halfway between the incident light direction and the reflection direction, $\hat{\bm{n}}$ is the macroscopic surface normal, and $\alpha=\hat{R}^{2}$ is the surface roughness parameter, ranging from [0,1]. As $\alpha$ approaches 0, the surface tends towards a mirror-like reflection, while $\alpha$ approaching 1 indicates an extremely rough surface.
@@
$F(\bm{\omega}_{o},\bm{h})=F_{0}+(1-F_{0})\cdot(1-(\bm{\omega}_{o}\cdot\bm{h}))^{5}$ (20)
where $F_{0}$ is the reflectance at normal incidence, also known as the base reflectance and $\hat{\bm{\omega}_{o}}\cdot\bm{h}$ is the dot product between the outgoing radiance and the half-vector, dictating the cosine of the angle of incidence. In our method, we assume dielectric materials with $F_{0}=0.04$ in the Fresnel term.
@@
Appendix 0.B Additional Implementation Details
0.B.1 Implemetaions of SIR
In Sec. 3.6, we use a three-phase training strategy to implement our framework, which includes the geometry and HDR-radiance field, irradiance and hard shadow field, and material and soft shadow estimation. The details are as follows: